OpenStudy (mony01):

Find volume of solid obtained by rotating region bounded by given curves about specified line. Sketch region, solid and disk or washer. y=sinx , y=cosx , 0

OpenStudy (dumbcow):

 \[V = \pi \int\limits_0^{\pi/4} (\cos x +1)^2 - (\sin x+1)^2 dx\]